# A Novel Method to Constrain Tidal Quality Factor from A Non-synchronized Exoplanetary System

### v1

@@ 4 Constraint of $Q^{\prime}$ in Target System
$=8.266+\left[-0.043-\log\left(\frac{\omega_{\mathrm{tide}}/2\pi}{2.48\,\mathrm{d^{-1}}}\right)\right]q.$
@@
Figure 6 also plots previous theoretical results based on the tidal interaction by the wave breaking mechanism (the green square; Barker & Ogilvie, 2010), the non-linear damping mechanism (the blue triangle; Essick & Weinberg, 2016), and the resonance locking (the red circle; Ma & Fuller, 2021) (see Appendix C for detail). The tidal quality factors derived from the wave braking and non-linear damping mechanisms are more than a hundred times smaller than our lower bound; in other words, these mechanisms predict considerably strong tidal torque, compared with that constrained by the NGTS-10 system. This suggests that the tidal efficiencies of these mechanisms should be weakened in the system with a slow rotator and a close-in gaseous planet. On the other hand, the resonance locking mechanism appears to provide a (marginally) consistent result with our constraint. However, Ma & Fuller (2021) show that this mechanism may be unlikely to be effective in the system of a low-mass star and a Jovian planet (see also Appendix C). At the moment, we do not know any plausible mechanism that could work in the non-synchronized system (see Section 5.2).
@@
$t_{\mathrm{decay}}$ $=\frac{4Q^{\prime}}{27}\frac{1}{\Omega_{\mathrm{orb}}}\left(\frac{M_{\mathrm{s}}}{M_{\mathrm{p}}}\right)\left(\frac{G(M_{s}+M_{p})}{\Omega_{\mathrm{orb}}^{2}R_{s}^{3}}\right)^{5/3}$
$\gtrsim 5.6\times 10^{-0.043q}\,\mathrm{Gyr}.$ (14)
@@ 5.1.4 Beyond Two-body system
The fundamental presumption of this paper is that we consider the isolated two-body system consisting of a star and a planet (see beginning of Section 2). However, when a third body is present, the evolution of the system will be altered by its gravity; a possible outcome of such three-body interaction is the high-eccentricity migration of a planet in the late evolutionary stage (e.g., Perryman, 2018, Section 10.12). If the NGTS-10 system had undergone such drastic migration recently, it could be transiently located in the forbidden region of the $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane; in this case, we cannot rule out strong tidal torque in the NGTS-10 system.
However, the probability of observing such a system is considered to be low because the timescale of orbital decay, $t_{\mathrm{decay}}$ (equation 14), is generally short in the presence of the strong tidal interaction. If we apply $Q^{\prime}=10^{6}$ as a moderately strong tidal torque to the NGTS-10 system, we obtain $t_{\mathrm{decay}}\approx 30\,\mathrm{Myr}$ from equation (14). This timescale is a tiny fraction of the inferred age of NGTS-10 $\gtrsim 7\,\mathrm{Gyr}$ (McCormac et al., 2020); the probability of a recent three-body interaction in the NGTS-10 system would not be high.
@@ 5.2 Implication from obtained weak tide
Let us begin the discussion with the case of positive $q$, described by redder lines in Figure 7. In this case, the tidal torque is suppressed as a system approaches co-rotation ($\omega_{\mathrm{tide}}=0\,\mathrm{d^{-1}}$). Then, the tidal interaction in a synchronized system, which satisfies $Q^{\prime}\gtrsim 10^{8.5}$, is weak enough to be negligible. This tidal properties cannot explain observed synchronized systems with low eccentricity and low inclination (e.g., Albrecht et al., 2012, see also Section 1). In order to solve this inconsistency, we should consider alternative tidal mechanisms, which work for a synchronized system effectively, in addition to the tidal interaction with $q>0$ constrained by the NGTS-10 system. For example, the dynamical tide through the inertial wave is considered to be a reliable mechanism (e.g., Mathis, 2015).
On the other hand, the case of negative $q$, which is described by bluer lines in Figure 7, expects a strong tidal interaction in a synchronized system. Therefore, it is not necessary to consider additional effects in this case. Examples of tidal mechanism with negative $q$ include the resonance locking, which satisfies our constraint (Figure 6).
Long-term observations of the shift in transit timing (e.g., Ivshina & Winn, 2022) are a powerful tool to pin down the constraint on the tidal interaction working in synchronized systems. Theoretical investigation with self-consistent numerical calculations (e.g., Ogilvie, 2014) is also essential to elucidate the dominant tidal process operating in various types of two-body systems.
@@ 5.3 Extension to eccentric and misaligned systems
Our analysis has been restricted on planets in circular orbits on the co-planar plane with the stellar spin. An extension to close-in planets with finite eccentricity, $e$, and inclination angle between the orbital axis and stellar spin axis, $i$, is meaningful because planets with large $e$ or $i$ have been identified with rate of 5-15 percent (Winn et al., 2010; Albrecht et al., 2012, 2022; Morgan et al., 2024).By employing formulations developed by previous works (Hut, 1981; Eggleton et al., 1998; Ferraz-Mello et al., 2008; Barker & Ogilvie, 2009; Xue et al., 2014), we can formulate AM co-evolution of eccentric and misaligned planets. We expect to obtain a forbidden region formed by a three dimensional unstable manifold in the four-dimensional phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}},i,e$). We note that our results in the current paper is a projection of such general cases onto the two dimensional plane with $i=0$ and $e=0$.
@@ 5.4 Applications to different objects
A promising candidate for such massive companions that readily comes to mind is a brown dwarf. An additional advantage of a brown dwarf around a low-mass star is that its formation is supposed to proceed simultaneously with that of the host star (see, e.g., Offner et al., 2023), which enables us to apply the initial conditions for planetary systems (Section 2.2) to brown-dwarf systems with a simple extrapolation. We expect the non-synchronized system with a low-mass star and a brown dwarf to provide a tight constraint of $Q^{\prime}$ even if $\Omega_{\mathrm{orb}}/\Omega_{\mathrm{spin}}(>1)$ is relatively small. A caveat is that such systems may be rare firstly because the number of close-in brown dwarfs orbiting K-type stars is depleted compared to brown dwarfs orbiting other types of stars (Grieves et al., 2021) and secondly because a system with a close-in brown dwarf and an M-type star is easily synchronized owing to their small host-to-companion mass ratio (Fukui et al. in prep).
@@ 6 Summary and Conclusion
Our framework can provide a tight constraint on tidal interaction in a system composed of a slow rotator and a close-in Jovian planet. We demonstrated that NGTS-10 – NGTS-10 b, the system with the currently known largest $\Omega_{\mathrm{orb}}/\Omega_{\mathrm{spin}}$, only allows extremely weak tide , $Q^{\prime}\gtrsim 10^{8}$. None of the currently known theoretical models can explain such weak tidal interaction. Additionally, the predicted shift of transit timing due to the tidal interaction for this system is too small to detect with the current observational accuracy within a few decades.
@@ Acknowledgement
T.T. is supported by IGPEES, WINGS Program in the University of Tokyo. T.K.S. is supported by Grants-in-Aid for Scientific Research from the MEXT/JSPS of Japan, 22H01263.
@@ Appendix D Calculation of $I_{\mathrm{s}}$
Stellar abundance Solar metallicity 1
@@
We determine $I_{\mathrm{s}}$, which is assumed to be time-independent (Section 2), by using mesa stellar evolution code (r12278) with the input parameters summarized in Table 5131313This setting reproduces the physical properties of the present Sun well.. To follow the evolution of NGTS-10, we calculate the time evolution of a $M_{\mathrm{s}}=0.7M_{\odot}$ star from the ZAMS to 14 Gyr ($\approx$ the age of the universe), which is before the end of the MS phase. Figure 10 presents the time variation of a dimensionless quantity, gyration radius, $\hat{r}_{\mathrm{g}}\equiv\sqrt{I_{\mathrm{s}}/(M_{\mathrm{s}}R_{\mathrm{s}}^{2})}$ (cf. Damiani & Lanza, 2015; Damiani & Díaz, 2016). $\hat{R}_{\mathrm{g}}$ slowly decreases with time as the star evolves to core-halo-like structure. We are employing
@@ Appendix E Autonomous system
The theoretical background of this paper is based on the fact that the systems consisting of equations (3) and (4) can be treated as two-dimensional autonomous systems under some presumptions. In this section, we briefly introduce the proof of the property of the autonomous system used in Section 2.2 (e.g., Arnold, 2006; Hirsch et al., 2012, for detail).
@@
where $t$, $x_{i}(t)$, and $f_{i}$ (for $i=1,\cdots,N$) represent the time as the independent variable, the element of the Euclid space as a dependent variable, and the function whose arguments are $x_{i}$, respectively. For abbreviation, we use the vector notation $\bm{x}=(x_{1},\cdots,x_{n})^{\mathrm{T}}$ and $\bm{f}(\bm{x})=(f_{1}(\bm{x}),\cdots,f_{n}(\bm{x}))^{\mathrm{T}}$, where “T” means transposition of a vector. At this time, the solution of equation (E1) is shown as the trajectory in the $n$-dimensional phase space of $(x_{1},\cdots,x_{n})$. Then, we can deduce that the tangent vector of the trajectory is equal to $\bm{f}(\bm{x})$.
Here, let $\bm{x_{*}}(t)$ be the unique solution of the initial value problem for equation (E1) when $\bm{x}(0)=\bm{x}_{\mathrm{init}}$, where $\bm{x}_{\mathrm{init}}$ is a particular vector. In this case, $\bm{x_{*}}(t-t_{0})$ for an arbitrarily selected time $t_{0}$ solves the initial value problem for equation (E1) when $\bm{x}(t_{0})=\bm{x}_{\mathrm{init}}$. Therefore, we find that the trajectory is determined independently of the value of $t$. Moreover, it indicates that two trajectories that pass through the same point are identical, which is equivalent to that two different trajectories have no intersection.

### v2

@@ 4 Constraint of $Q^{\prime}$ in Target System
$=8.268+\left[-0.045-\log\left(\frac{\omega_{\mathrm{tide}}/2\pi}{2.48\,\mathrm{d^{-1}}}\right)\right]q.$
@@
Figure 6 also plots previous theoretical results based on the tidal interaction by the wave breaking mechanism (the green square; Barker & Ogilvie, 2010), the non-linear damping mechanism (the blue triangle; Essick & Weinberg, 2016), and the resonance locking (the red circle; Ma & Fuller, 2021) (see Appendix C for detail). The tidal quality factors derived from the wave braking and non-linear damping mechanisms are more than a hundred times smaller than our lower bound; in other words, these mechanisms predict considerably strong tidal torque, compared with that constrained by the NGTS-10 system. This suggests that the tidal efficiencies of these mechanisms should be weakened in the system with a slow rotator and a close-in gaseous planet. On the other hand, the resonance locking mechanism appears to provide a consistent result with our constraint as this mechanism can quantitatively explain the current position of the NGTS-10 system. However, Ma & Fuller (2021) show that this mechanism may be unlikely to be effective in the system of a low-mass star and a Jovian planet (see also Appendix C). At the moment, we do not know any plausible explanation about the weak tide of this system (see Section 5.2 for detail).
@@
$t_{\mathrm{decay}}$ $=\frac{4Q^{\prime}}{27}\frac{1}{\Omega_{\mathrm{orb}}}\left(\frac{M_{\mathrm{s}}}{M_{\mathrm{p}}}\right)\left[\frac{G(M_{s}+M_{p})}{\Omega_{\mathrm{orb}}^{2}R_{s}^{3}}\right]^{5/3}$
$\gtrsim 5.6\times 10^{-0.045q}\,\mathrm{Gyr}.$ (14)
@@ 5.1.4 Beyond Two-body system
The fundamental presumption of this paper is that we consider the isolated two-body system consisting of a star and a planet (see beginning of Section 2). However, when a third body is present, the evolution of the system will be altered by its gravity; a possible outcome of such three-body interaction is the high-eccentricity migration of a planet in the late evolutionary stage (cf. Masuda, 2017). If the NGTS-10 system had undergone such drastic migration recently, it could be transiently located in the forbidden region of the $\Omega_{\mathrm{orb}}$$\Omega_{\mathrm{spin}}$ plane; in this case, we cannot rule out strong tidal torque in the NGTS-10 system.
However, the probability of observing such a system is considered to be low because the timescale of orbital decay, $t_{\mathrm{decay}}$ (equation 14), is generally short in the presence of the strong tidal interaction. If we apply $Q^{\prime}=10^{6}$ as a moderately strong tidal torque to the NGTS-10 system, we obtain $t_{\mathrm{decay}}\approx 30\,\mathrm{Myr}$ from equation (14). This timescale is a tiny fraction of the inferred age of NGTS-10 order than $7\,\mathrm{Gyr}$ (McCormac et al., 2020); the probability of a recent three-body interaction in the NGTS-10 system would not be high.
@@ 5.2 Implication from obtained weak tide
Let us begin the discussion with the case of positive $q$, described by redder lines in Figure 7. In this case, the tidal torque is suppressed as a system approaches co-rotation ($\omega_{\mathrm{tide}}=0\,\mathrm{d^{-1}}$). Then, the tidal interaction in a synchronized system, which satisfies $Q^{\prime}\gtrsim 10^{8.5}$, is weak enough to be negligible. This tidal properties cannot explain observed synchronized systems with low inclination (e.g., Albrecht et al., 2012, 2022, see also Section 1). In order to solve this inconsistency, we should consider alternative tidal mechanisms, which work for a synchronized system effectively, in addition to the weak tidal interaction with $q>0$ constrained by the NGTS-10 system. For example, the dynamical tide through the inertial wave is considered to be a possible mechanism (e.g., Mathis, 2015).
On the other hand, the case of negative $q$, which is described by bluer lines in Figure 7, expects a strong tidal interaction in a synchronized system. Therefore, it is not necessary to consider additional effects in this case. Examples of tidal mechanism with negative $q$ include the resonance locking, which satisfies our constraint (Figure 6). Therefore, it can be a plausible tidal mechanism if further theoretical examination shows that the resonance locking mechanism is also valid for Jovian planets (see also Appendix C).
Long-term observations of the shift in transit timing (e.g., Ivshina & Winn, 2022; Adams et al., 2024) are a powerful tool to pin down the constraint on the tidal interaction working in synchronized systems. Theoretical investigation with self-consistent numerical calculations (e.g., Ogilvie, 2014) is also essential to elucidate the dominant tidal process operating in various types of two-body systems.
@@ 5.3 Extension to eccentric and misaligned systems
Our analysis has been restricted on planets in circular orbits on the co-planar plane with the stellar spin. An extension to close-in planets with finite eccentricity, $e$, and spin-orbit inclination angle, $i$, is meaningful because planets with large $e$ or $i$ have been identified with rate of 5-15 percent (Winn et al., 2010; Albrecht et al., 2012, 2022; Morgan et al., 2024). By employing formulations developed by previous works (Hut, 1981; Eggleton et al., 1998; Ferraz-Mello et al., 2008; Barker & Ogilvie, 2009; Xue et al., 2014), we can formulate AM co-evolution of eccentric and misaligned planets. We expect to obtain a forbidden region formed by a three dimensional unstable manifold in the four-dimensional phase space of ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}},e,i$). In such a case, we should consider the tidal dissipation not only in the stellar interior but also in the planetary interior, both of which affect the shape of trajectories in the ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}},e,i$) space (Jackson et al., 2008; Bonomo et al., 2017; Efroimsky & Makarov, 2022). Therefore, the shape of the unstable manifolds with respect to the $e$ and $i$ directions is considered to depend on the efficiencies of these two tidal processes. A quantitative investigation of these dependencies is a subject for future work.
It should be noted that the unstable manifolds used in the current paper correspond to the projections of such general cases onto the two-dimensional plane at $e=0$ and $i=0$. Unstable manifolds for finite $e$ and $i$ projected onto that plane are different from the unstable manifold for $e=i=0$. However, from the characteristic properties of the unstable manifold, the system evolves while ensuring that ($\Omega_{\mathrm{orb}},\Omega_{\mathrm{spin}}$) remain outside the forbidden region determined by $(e,i)$ at any given time. Therefore, provided that the system has $e=i=0$ at the time of observation, we can safely use the forbidden region in Figure 2 to constrain the tidal torque, regardless of the evolutionary history of $e$ and $i$ (see Section 2.3).
@@ 5.4 Applications to different objects
A promising candidate for such massive companions that readily comes to mind is a brown dwarf. An additional advantage of a brown dwarf around a low-mass star is that its formation is supposed to proceed simultaneously with that of the host star (see, e.g., Offner et al., 2023), which enables us to apply the initial conditions for planetary systems (Section 2.2) to brown-dwarf systems with a simple extrapolation. We expect the non-synchronized system with a low-mass star and a brown dwarf to provide a tight constraint of $Q^{\prime}$ even if $\Omega_{\mathrm{orb}}/\Omega_{\mathrm{spin}}(>1)$ is relatively small. A caveat is that such systems may be rare firstly because the number of close-in brown dwarfs orbiting K-type stars is depleted compared to brown dwarfs orbiting other types of stars (Grieves et al., 2021) and secondly because a system with a close-in brown dwarf and an M-type star is easily synchronized owing to their small host-to-companion mass ratio.
@@ 6 Summary and Conclusion
Our framework can provide a tight constraint on tidal interaction in a system composed of a slow rotator and a close-in Jovian planet. We demonstrated that NGTS-10 – NGTS-10 b, the system with the currently known largest $\Omega_{\mathrm{orb}}/\Omega_{\mathrm{spin}}$, only allows extremely weak tide , $Q^{\prime}\gtrsim 10^{8}$. This constraint is in conflict with the $Q^{\prime}$ values predicted by the wave braking mechanism or the nonlinear-damping mechanism. Among the currently known mechanisms, only the resonance locking is consistent with this constraint, although this mechanism may not be effectively working in Jovian planets (Ma & Fuller, 2021). Additionally, our constraint predicts that the shift of transit timing due to the tidal interaction for this system is too small to detect with the current observational accuracy within a few decades.
@@ Acknowledgement
T.T. is supported by IGPEES, WINGS Program in the University of Tokyo and Research Fellowships for Young Scientists in JSPS (24KJ0605). T.K.S. is supported by Grants-in-Aid for Scientific Research from the MEXT/JSPS of Japan, 22H01263.
@@ Appendix D Calculation of $I_{\mathrm{s}}$
Initial abundance The solar metallicity 1
@@
We determine $I_{\mathrm{s}}$, which is assumed to be time-independent (Section 2), by using mesa stellar evolution code (r12278) with the input parameters summarized in Table 5131313This setting reproduces the physical properties of the present Sun well. We upload the inlists in Zenodo: https://doi.org/10.5281/zenodo.11441698 (catalog 10.5281/zenodo.8385153) . To follow the evolution of NGTS-10, we calculate the time evolution of a $M_{\mathrm{s}}=0.7M_{\odot}$ star from the ZAMS to 14 Gyr ($\approx$ the age of the universe), which is before the end of the MS phase. Figure 10 presents the time variation of a dimensionless quantity, gyration radius, $\hat{r}_{\mathrm{g}}\equiv\sqrt{I_{\mathrm{s}}/(M_{\mathrm{s}}R_{\mathrm{s}}^{2})}$ (cf. Damiani & Lanza, 2015; Damiani & Díaz, 2016). $\hat{R}_{\mathrm{g}}$ slowly decreases with time as the star evolves to core-halo-like structure. We are employing
@@ Appendix E Autonomous system
The theoretical background of this paper is based on the fact that the systems consisting of equations (3) and (4) can be treated as two-dimensional autonomous systems under some presumptions. In this section, we briefly introduce the proof of the property of the autonomous system used in Section 2.2 (e.g., Arnold, 1992; Hsieh & Sibuya, 1999, for detail).
@@
where $t$, $x_{i}(t)$, and $f_{i}$ (for $i=1,\cdots,N$) represent the time as the independent variable, the element of the Euclid space as a dependent variable, and the Lipschitz-continous function whose arguments are $x_{i}$, respectively. For abbreviation, we use the vector notation $\bm{x}=(x_{1},\cdots,x_{n})^{\mathrm{T}}$ and $\bm{f}(\bm{x})=(f_{1}(\bm{x}),\cdots,f_{n}(\bm{x}))^{\mathrm{T}}$, where “T” means transposition of a vector. At this time, the solution of equation (E1) is shown as the trajectory in the $n$-dimensional phase space of $(x_{1},\cdots,x_{n})$. Then, we can deduce that the tangent vector of the trajectory is equal to $\bm{f}(\bm{x})$.
Here, let $\bm{x_{*}}(t)$ be the unique solution of the initial-value problem for equation (E1) when $\bm{x}(0)=\bm{x}_{\mathrm{init}}$, where $\bm{x}_{\mathrm{init}}$ is a particular vector. In this case, $\bm{x_{*}}(t-t_{0})$ for an arbitrarily selected time $t_{0}$ solves the initial-value problem for equation (E1) when $\bm{x}(t_{0})=\bm{x}_{\mathrm{init}}$. Therefore, we find that the trajectory in the phase space is determined independently of the value of $t$. Moreover, it indicates that two trajectories that pass through the same point are identical, which is equivalent to that two different trajectories have no intersection.